OpenStudy (anonymous):

What is the area of square whose diagonal is 10 cm?

hartnn (hartnn):

can u find side using pythagoras theorem ?

OpenStudy (anonymous):

40?

OpenStudy (jiteshmeghwal9):

\[\huge{\frac{1}{2} of (d)^2}\]\[\huge{\frac{1}{2} \times (10)^2=?}\]

hartnn (hartnn):

nopes, what u got side as ?

OpenStudy (jiteshmeghwal9):

now slve this is correct :)

OpenStudy (jiteshmeghwal9):

@Faye281

OpenStudy (anonymous):

50?>

hartnn (hartnn):

yes

OpenStudy (jiteshmeghwal9):

ya !

OpenStudy (anonymous):

thanks :))

OpenStudy (jiteshmeghwal9):

yw:)